OpenStudy (anonymous):

Express 2x^2 +9x -6/(x^2 + x -6) as a partial fraction

OpenStudy (anonymous):

divide first

OpenStudy (anonymous):

Dont we split the bottom into the components? factoring the bottom to (x+3)(x-2) ?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

the numerator and denominator are both polynomials of degree 2, so first you have to divide you will get 2 + something and the "something" is what you will express as partial fractions

OpenStudy (anonymous):

Ahhh i see, this is part 2 of my investigation. Algebraic long division. Okie doke i'll do that first.

OpenStudy (anonymous):

So doing that we end up with 2 + 7x+6/x^2+x-6 ?

OpenStudy (anonymous):

this problem has been cooked up to be easy you will get nice integers for the partial fractions

OpenStudy (anonymous):

right

OpenStudy (anonymous):

if the numerator has degree greater than or equal to the denominator, you need to divide first, get a quotient and a remainder, then you can find partial fractions for the remainder kind of like if you had \(\frac{24}{10}\) you might write is as \(2\tfrac{4}{10}\) and then mess around with the \(\frac{4}{10}\) part

OpenStudy (anonymous):

Thankyou very much, i think i've got it now! I've got the answer as 2+ (3/x+3) + (4/x-2)

OpenStudy (anonymous):

I'd never come across a question like that before, and i've got a test on it tomorrow :\ Hopefully i'll do ok :S

OpenStudy (anonymous):

i think so there is a very snappy way to do it, do you know how?

OpenStudy (anonymous):

No i dont? I would like to see it if you have time?

OpenStudy (anonymous):

forgetting about the 2, and looking only at \[\frac{7x+6}{(x+3)(x-2)}\] you know this is going to be \[\frac{A}{x+3}+\frac{B}{x-2}\] so to find B, put your finger over the \(x-2\) and the denominator \[\frac{7x+6}{(x+3)\cancel{(x-2)}}\]and then replace \(x\) by 2 to get \(\frac{14+6}{5}=4\)

OpenStudy (anonymous):

then to find A put your finger over the \(x+3\) in the denominator \[\frac{7x+6}{\cancel{(x+3)}(x-2)}\] and replace \(x\) by \(-3\) to get \[\frac{-21+6}{-5}=3\]

OpenStudy (anonymous):

Oh wow.... does this work with every equation?

OpenStudy (anonymous):

well it works for ones where the denominator has non repeated linear factors

OpenStudy (anonymous):

so i wouldn't work in the case of say \(x^2+x+1\) in the denominator, or \((x-4)^2\)

OpenStudy (anonymous):

it is really just a gimmick, it works exactly the same way as if you had written \[A(x-2)+B(x+3)=7x+6\] and then replaced \(x\) by \(2\) just solves in one step is all

OpenStudy (anonymous):

Hmm, would you advise doing it the long way in a test? because i have a test on this tomorrow :\

OpenStudy (anonymous):

whatever you are more comfortable with it really only saves a second from writing \[A(x-2)+B(x+3)=7x+6\] and then putting \(x=2\) to get \(5B=7\times 2+6\)

OpenStudy (anonymous):

the most common mistake is to forget which one you are solving for, so make sure on your paper to write \[\frac{A}{x+3}+\frac{B}{x-2}\] so you an remember which one is over each factor

OpenStudy (anonymous):

Hmmm i've never learnt that before. I went with ax-2a+bx+3b = 7x + 6 then x(a+b) -2a +3b (a+b)=7 and 3b-2a=6 then solve those two simultaneously

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

now you have to solve a system of equations

OpenStudy (anonymous):

apparently you did not have much trouble doing it, but it seems to me the long way to do things, and it is almost always unnecessary

OpenStudy (anonymous):

so we can just plug x=2 into the a term to cancel the a and then solve? Does this always work? assuming one term is cancelled?

OpenStudy (anonymous):

multiply out collect terms equate like coefficients solve the system you might need this when you do laplace transforms, but for algebra it is too much work

OpenStudy (anonymous):

No this isnt in our syllabus. Im in the highest highschool maths that is non-specialist. but this isnt in our course. They call it an "extended piece of work" or EPW which is essentially a mathematical investigation.

OpenStudy (anonymous):

yeah you can almost always use this method

OpenStudy (anonymous):

i think i recall that my math teacher was an "extended piece of work"

OpenStudy (anonymous):

Hahahaha mine is great, we have a class of 6 (myself included)

OpenStudy (anonymous):

here is an example, the answer will not be so nice but the method is easy \[\frac{2x+3}{(x-4)(x+3)}\]

OpenStudy (anonymous):

you know it is \[\frac{A}{x-4}+\frac{B}{x+3}\] so you know that \[A(x+3)+B(x-4)=2x+3\] and the point is that this has to be true NO MATTER WHAT X IS

OpenStudy (anonymous):

therefore you are free to pick any \(x\) you like and the obvious choice is \(x=4\) because that will get rid of the B term and leave you with \[7A=2\times 4+3\] or \[7A=11\] so \(A=\frac{11}{7}\) much easier than multiplying out, collecting terms and then solving a system

OpenStudy (anonymous):

but do whatever you are comfortable with

OpenStudy (anonymous):

So just testing my skill. sub-ing x=-3 gives b=3/7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

easy right? just make sure to recall which one is A and which one is B, in other words, don't skip the step of writing \(\frac{A}{x-4}+\frac{B}{x+3}\) somewhere on your paper so you can refer back to it

OpenStudy (anonymous):

good luck on your test

OpenStudy (anonymous):

@satellite73 sorry mate, i lost connection then. But Thankyou very much for all of the help! I definitely owe you one. And so I just wanted to wish you all the best, and a happy and safe day/night, wherever you are. Its night here in aus. Thankyou again for all the help mate, you're a bloody legend!

OpenStudy (anonymous):

you are quite welcome. i am still laughing at "piece of work" maybe they don't have that idiom in aus

OpenStudy (anonymous):

@satellite73 Hahaha no we do :P It was pretty funny. Just a quick question. How do you know which term goes first, so you dont solve the a and b wrong? so if you have (x-3)(x+4) how do you know which term to make a and which to make b? Im assuming a wrong choice would reverse the values?